# Modal logics of almost sure validities in some classes of euclidean and transitive frames

### v1

@@ 4 Transitive frames
Given a transitive Noetherian relation $R$ on a set $X,$ the transitive reduction of $R$ is defined as $R^{-}=R\setminus R^{2}.$ It is the smallest relation on $X$ such that ${R\subseteq(R^{-})^{*}}.$
If $F=(X,\,R)$ is a finite connected frame and $F\models\mathbf{GL.3}$ or $F\models\mathbf{Grz.3}$ then $(X,\,R^{-})$ is an inverse tree: there is a unique $R^{-}$-maximal element $a_{0}$, and for any $a\in X,$ $|R^{-}(a)|\leq 1$ and $a(R^{-})^{*}a_{0}.$
Note that if $(X,\,R)\in\operatorname{Con}\operatorname{Fr}\mathbf{GL.3}$ then $R=(R^{-})^{+}$. Similarly, if $(X,\,R)$ is a frame in $\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3}$ then $R=(R^{-})^{*}.$ Therefore the mapping $R\mapsto R^{-}$ has an inverse on $\operatorname{Con}\operatorname{Fr}\mathbf{GL.3}$ and on $\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3}.$
@@
For any $n\in\omega,$ the following sets are bijective:
[…]
    $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{GL.3};$
[…]
    $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3};$
[…]
    the set $T_{n}$ of inverse trees over the states $n.$
Moreover, the bijection preserves the distance $d_{R}(a,\,b)=\min\{n\in\omega:\>aR^{n}b\}$ for any $a,\,b\in n.$
Proof 4.2.
@@
$\lim_{n\to\varnothing}\operatorname{P}\left(\exists a\in n:\>|\hat{R}_{n}(a)|\geq r\right)=\lim_{n\to\infty}\operatorname{P}(\hat{h}_{n}>r)\geq q.$
With an asymptotic probability $q>0$ there exists $a\in n$ such that ${|\hat{R}_{n}(a)|\geq r}.$ In this case $\hat{F}_{n}(\mathbf{GL.3}){\uparrow}a$ is an irreflexive chain of cardinality at least $r$, so $F$ is isomorphic to some generated subframe of $\hat{F}_{n}(\mathbf{GL.3}){\uparrow}a.$
@@
Therefore $\mathbf{GL.3}^{\mathbf{as}}\subseteq\mathbf{GL.3}$ by Theorem 2.7. The converse is true by Theorem 1.1.
@@ 5 Results and discussion
$\mathbf{K5}^{\mathbf{as}}=\mathbf{KD5}^{\mathbf{as}}=\mathbf{KD5};$ $\mathbf{K45}^{\mathbf{as}}=\mathbf{KD45}^{\mathbf{as}}=\mathbf{KD45};$ $\mathbf{K5B}^{\mathbf{as}}=\mathbf{K5B};$
@@
Our results on $\mathbf{KD5}^{\mathbf{as}},\,\mathbf{KD45}^{\mathbf{as}},\,\mathbf{K5B}^{\mathbf{as}},\,\mathbf{S5}^{\mathbf{as}},\,\mathbf{Grz.3}^{\mathbf{as}},\,\mathbf{GL.3}^{\mathbf{as}}$ provide examples of logics that are equal to their ‘almost sure’ counterparts. Finding a criterion that characterizes the logics with this property is an interesting direction for future research.

### v2

@@ 4 Transitive frames
Definition 4.1.
A frame $F=(X,\,R)$ is an inverse tree if there is a unique element $a_{0}$ such that $R^{-}(a_{0})=\varnothing$, and for any $a\in X\setminus\{a_{0}\},$ $|R^{-}(a)|=1$ and $a(R^{-})^{*}a_{0}.$ We denote $T_{n}$ the set of all inverse trees over the set of states $[n].$
Definition 4.2.
The distance between states $a$ and $b$ of a frame $F=(X,\,R)$ is the number $d_{R}(a,\,b)=\min\{n\in\omega:\>aR^{n}b\}$ for any $a,\,b\in[n].$
Definition 4.3.
Definition 4.4.
Given a transitive Noetherian relation $R$ on a set $X,$ the transitive reduction of $R$ is defined as $R^{-}=R\setminus R^{2}.$
It is straightforward to see that for a transitive Noetherian relation $R,$ its transitive reduction $R^{-}$ is the smallest relation on $X$ such that ${R\subseteq(R^{-})^{*}}$ [AGU72]. If $(X,\,R)\in\operatorname{Con}\operatorname{Fr}\mathbf{GL.3}$, then $R=(R^{-})^{+}$. Similarly, if $(X,\,R)$ is a frame in $\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3}$, then $R=(R^{-})^{*}.$ Therefore the mapping $R\mapsto R^{-}$ has an inverse on $\operatorname{Con}\operatorname{Fr}\mathbf{GL.3}$ and on $\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3}.$
@@
For any $n\in\omega,$ the sets of frames $T_{n},$ $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{GL.3},$ and $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{Grz.3},$ are bijective. Moreover, the bijection preserves the distance between any pair of states.
Proof 4.6.
@@
$\lim_{n\to\infty}\operatorname{P}\left(\exists a\in[n]:\>|\hat{R}_{n}(a)|\geq r\right)=\lim_{n\to\infty}\operatorname{P}(\hat{h}_{n}>r)\geq q.$
With an asymptotic probability $q>0$ there exists $a\in[n]$ such that ${|\hat{R}_{n}(a)|\geq r}.$ In this case $\hat{F}_{n}(\mathbf{GL.3}){\uparrow}a$ is an irreflexive chain of cardinality at least $r$, so $F$ is isomorphic to some generated subframe of $\hat{F}_{n}(\mathbf{GL.3}){\uparrow}a.$
@@
Since $([n],<)$ is a connected $\mathbf{GL.3}$-frame, $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{GL.3}\neq\varnothing$ for all $n\in\omega$, thus by Theorem 2.9 $\mathbf{GL.3}^{\mathbf{as}}\subseteq\mathbf{GL.3}$. The converse is true by Theorem 1.1.
@@ 5 Results and discussion
We developed several general results about the almost sure validities in random Kripke frames. Theorem 1.1 states that the almost sure validities in the random frame $\hat{F}(\mathcal{C})$ is a normal modal logic that extends $\mathcal{C}$ for any class of frames $\mathcal{C}$. Theorem 2.9, which applies for many commonly studied modal logics, establishes an important inclusion $L^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}L)$.
@@
$\mathbf{KD5}^{\mathbf{as}}=\mathbf{KD5};$ $\mathbf{KD45}^{\mathbf{as}}=\mathbf{KD45};$ $\mathbf{K5B}^{\mathbf{as}}=\mathbf{K5B};$
@@
Our results on $\mathbf{KD5},\,\mathbf{KD45},\,\mathbf{K5B},\,\mathbf{S5},\,\mathbf{Grz.3},\,\mathbf{GL.3}$ provide examples of logics that are equal to their ‘almost sure’ counterparts. Finding a criterion that characterizes the logics with this property is an interesting direction for future research.